The sum of two numbers is 24. the sum of their square is 338. Find the numbers.
Answer by stanbon(75887)

The sum of two numbers is 24. the sum of their square is 338. Find the numbers.
--------
x + y = 24
x^2 + y^2 = 338
-----
Substitute for "x" and solve for "y"::
(24-y)^2 + y^2 = 338
-----
24^2 - 48y + 2y^2 = 338
2y^2 - 48y + 238 = 0
----
y^2 - 24y + 119 = 0
-----
Factor:
(y-7)(y-17) = 0
----
Ans: The numbers are 7 and 17
---------------
Cheers,
Stan H.
---------------